OpenStudy (anonymous):

What is the common ratio of the sequence 80, 20, 5, 5/4 . . . .? What are the next three terms of the sequence?

OpenStudy (anonymous):

you are dividing by 4 in each sequence... So... (5/4) / 4 = 5/16 (5/16) /4 = 5/64 (5/64) /4 = 5/256

OpenStudy (anonymous):

Luis, you had the first two in the sequence correct, however the last one is 5/64 / 4 which is essentially 5/(4*64) = 5/256